# Proxy-small objects present compactly generated categories

### v1

@@ 1. Introduction
Proxy-smallness is a relatively new concept that was introduced in [5, 6] as a weakening of smallness, also called compactness, in triangulated categories. It is general enough that the residue field of any noetherian local ring is proxy-small in the derived category, but strong enough to yield useful formulas for local cohomology. The most striking application has been in commutative algebra. Dwyer, Greenlees, and Iyengar proved [5]*Theorem 9.4 that if a commutative noetherian ring $R$ is complete intersection then every object of $\mathsf{D}^{\mathrm{b}}(\operatorname{\mathsf{mod}}R)$ is proxy-small and they asked if this property characterized complete intersections. This remarkable statement turns out to be true, as was proved by Pollitz [16]. This opens many doors. By analogy it furnishes us with an elegant and homotopy invariant definition of complete intersection in derived settings, and it supports rather appealing, if speculative, connections to questions concerning finite generation of cohomology. The purpose of this article is to give a conceptual interpretation of proxy-smallness in analogy with the definition of compactness, that is in terms of some functor preserving coproducts.
Let $\mathsf{K}$ be an enhanced compactly generated triangulated category. Given a compact object $k\in\mathsf{K}$ there is an equivalence between the localizing subcategory $\operatorname{\mathrm{loc}}(k)$ generated by $k$ and the derived category $\mathsf{D}(E)$ where $E$ is the derived endomorphism ring $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(k,k)$ of $k$. This equivalence is given by the adjoint pair of functors $(-)\otimes_{E}k$ and $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(k,-)$, and boils down to the statements $E=\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(k,k)$ and $E\otimes_{E}k=k$, since both functors preserve coproducts. For a non-compact object $P$ this statement is false in general —the localizing subcategory $\operatorname{\mathrm{loc}}(P)$ depends on how $P$ sits in $\mathsf{K}$ in a subtle way, and not only on its endomorphism ring. For instance, in the derived category of abelian groups $\mathsf{D}(\mathbb{Z})$ the objects $\mathbb{Z}_{p}$ and $\mathbb{Z}[p^{-1}]/\mathbb{Z}$ have the same derived endomorphism ring but generate different localizing subcategories. The slogan is: giving a proxy-small object $P$ is the same thing as giving a presentation of the compactly generated localizing subcategory $\operatorname{\mathrm{loc}}(P)$ as a quotient of $\mathsf{D}(\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(P,P))$ (see 4.9).
@@ 2. Proxy-smallness revisited
Let us fix a compactly generated stable $\infty$-category (i.e. an $\aleph_{0}$-presentable stable $\infty$-category) with homotopy category $\mathsf{K}$. Essentially $\mathsf{K}$ is a compactly generated triangulated category (see [12]*Remark 1.4.4.3) and we can speak of mapping spaces, homotopy (co)limits, and so on. Nothing is lost if the reader prefers to think of instead fixing a differential graded, or model, category with compactly generated homotopy category. In fact, essentially all arguments take place at the level of homotopy categories. We denote the compact objects $\mathsf{K}^{\mathrm{c}}$ and recall that this category is, by definition, essentially small.
Throughout by ‘localization’ we will mean left Bousfield localization, i.e. a functor with a fully faithful right adjoint. For a set of objects $\mathsf{X}$ of $\mathsf{K}$ we denote by $\operatorname{thick}(\mathsf{X})$ and $\operatorname{\mathrm{loc}}(\mathsf{X})$ the smallest triangulated subcategories containing $\mathsf{X}$ and closed under summands and arbitrary coproducts respectively.
@@
Throughout we regard $\mathbf{p}$ as a full subcategory of $\mathsf{K}$. Using the enhancement we may consider the ind-completion $\mathsf{D}(\mathbf{p})$ of $\operatorname{thick}(\mathbf{p})$ and this identifies with the full subcategory $\operatorname{\mathrm{loc}}(\mathbf{p})$ of $\mathsf{K}$.
@@
In fact one could, and in some cases should, go further: we could also replace $P$ by a set of objects. Everything works just as well if one takes this extra step. In our primary cases of interest we can get away with a single object $P$ and so we stick to this setting—replacing $P$ by a subcategory is essentially cosmetic.
@@
For the rest of this section $P$ is a proxy-small object in $\mathsf{K}$ with a set of proxies $\mathbf{p}$. We are primarily concerned with understanding
$\Gamma=i_{*}i^{!}\cong\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(\mathbf{p},-)\otimes_{\mathbf{p}}\mathbf{p}.$
@@
Assume $P$ is proxy-small. The functor $(-)\otimes_{E}P\colon\mathsf{D}(E)\longrightarrow\operatorname{\mathrm{loc}}(P)$ is a localization. Moreover, it preserves products.
@@
Our starting point is the observation that $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(P,-)$ induces an equivalence from $\operatorname{thick}(P)$ to $\mathsf{D}(E)^{\mathrm{c}}$ with inverse $(-)\otimes_{E}P$. In particular, $\mathbf{p}\subseteq\operatorname{thick}(P)$ embeds fully faithfully into $\mathsf{D}(E)$ which induces a colimit and compact preserving embedding $\lambda\colon\operatorname{\mathrm{loc}}(\mathbf{p})\longrightarrow\mathsf{D}(E)$. While the localizing subcategory $\operatorname{\mathrm{loc}}(P)$ must be taken in $\mathsf{K}$, we note that $\operatorname{\mathrm{loc}}(\mathbf{p})$ is unambiguous here: $\mathbf{p}$ consists of objects that are compact in both $\mathsf{D}(E)$ and $\mathsf{K}$, and so it does not matter where we form the localizing subcategory $\mathbf{p}$ generates. By Brown representability $\lambda$ has a right adjoint $\rho$. Of course, this is overkill: the right adjoint is just the obvious restriction functor (up to identifying $\operatorname{\mathrm{loc}}(\mathbf{p})$ and $\mathsf{D}(\mathbf{p})$).
@@
Consider $R=k[x]/(x^{2})$. In this case, $E=k[\theta]$ with $\theta$ in degree $1$ and no higher homotopical structure (i.e. $E$ is formal). We have
@@
This extends to any artinian commutative local ring, and in fact it is more general still.
@@
Let $R$ be a coherent ring and $G$ a classical generator for $\mathsf{D}^{\mathrm{b}}(\operatorname{\mathsf{mod}}R)$. We note that such a $G$ does not exist for all rings, not even commutative ones. Then $G$ is certainly proxy-small: $\operatorname{thick}(G)=\mathsf{D}^{\mathrm{b}}(\operatorname{\mathsf{mod}}R)$ which contains all perfect complexes and it follows that $\operatorname{\mathrm{loc}}(G)=\mathsf{D}(R)$ which is generated by the perfect complexes.
Hence, letting $E$ denote the derived endomorphism ring of $G$, we see $\mathsf{D}(R)$ is a localization of $\mathsf{D}(E)$ and $\mathsf{D}(E)^{\mathrm{c}}\cong\mathsf{D}^{\mathrm{b}}(\operatorname{\mathsf{mod}}R)$. Thus $\mathsf{D}(E)\cong\mathsf{K}(\operatorname{\mathsf{Inj}}R)$ and this localization is the usual one up to equivalence.
@@
Using Theorem 2.9 we can give an efficient proof of effective constructibility as in [6]*Theorem 4.10.
@@
$-\otimes_{\mathbb{Z}_{p}}P\colon\mathsf{D}(\mathbb{Z}_{p})\longrightarrow\mathsf{D}(\mathbb{Z}).$
This is not surprising when we keep in mind Theorem 2.9 that tells us $\mathsf{D}(E)\longrightarrow\mathsf{K}$ factors via a localization (which has no reason to be the trivial one).
@@ 3. Detecting smashing localizations
We continue the notation of the previous sections, with the caveat that we no longer assume that $P$ is necessarily proxy-small. If $P$ is proxy-small then the functor $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(P,-)\otimes_{E}P$ preserves coproducts, even though $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(P,-)$ does not if $P$ is not compact. It is then natural to ask what we can deduce about a general $P$ with the property that $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(P,-)\otimes_{E}P$ preserves coproducts.
@@
Since $\mathsf{K}$ is compactly generated $\operatorname{\mathrm{loc}}(P)$ is well-generated [15, Theorem 4.4.9] and so by Brown representability [15, Theorem 8.3.3] there is for all $X$ a localization triangle
@@ 4. Characterization of proxy-smallness
One can also prove the theorem by checking the definition of proxy-smallness directly using that $\mathsf{J}=\operatorname{\mathrm{loc}}(\pi(E))$, because $\pi$ is a localization, and applying Lemma 4.1.
@@
Remark 4.10.
In fact the localization corresponding to a proxy-small object also preserves products. This was shown in 2.9.
When $\mathsf{K}$ has a compact generator each compactly generated localizing subcategory comes with a canonical presentation. Let $\mathbf{p}$ be a set of compact objects of $\mathsf{K}$ and set

### v2

@@ Abstract.
Key words and phrases:
compactly generated triangulated category, complete intersection, small object, proxy-small object
2010 Mathematics Subject Classification:
18G80 (primary); 13D09 (secondary)
@@ 1. Introduction
Proxy-smallness is a relatively new concept that was introduced in [6, 7] as a weakening of smallness, also called compactness, in triangulated categories. It is general enough that the residue field of any noetherian local ring is proxy-small in the derived category, but strong enough to yield useful formulas for local cohomology. This has had striking applications, both in homotopy theory and in commutative algebra. For instance, Dwyer, Greenlees, and Iyengar proved [6]*Theorem 9.4 that if a commutative noetherian ring $R$ is complete intersection then every object of $\mathsf{D}^{\mathrm{b}}(\operatorname{\mathsf{mod}}R)$ is proxy-small and they asked if this property characterized complete intersections. This remarkable statement turns out to be true, as was proved by Pollitz [20]. This opens many doors. By analogy it furnishes us with an elegant and homotopy invariant definition of complete intersection in derived settings, and it supports rather appealing, if speculative, connections to questions concerning finite generation of cohomology. The purpose of this article is to give a conceptual interpretation of proxy-smallness in analogy with the definition of compactness, that is in terms of some functor preserving coproducts.
Let $\mathsf{K}$ be an enhanced compactly generated triangulated category. Given a compact object $k\in\mathsf{K}$ there is an equivalence between the localizing subcategory $\operatorname{\mathrm{loc}}(k)$ generated by $k$ and the derived category $\mathsf{D}(E)$ where $E$ is the derived endomorphism ring $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(k,k)$ of $k$. This equivalence is given by the adjoint pair of functors $(-)\otimes_{E}k$ and $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(k,-)$, and boils down to the statements $E=\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(k,k)$ and $E\otimes_{E}k=k$, since both functors preserve coproducts. For a non-compact object $P$ this statement is false in general —the localizing subcategory $\operatorname{\mathrm{loc}}(P)$ depends on how $P$ sits in $\mathsf{K}$ in a subtle way, and not only on its endomorphism ring. For instance, in the derived category of abelian groups $\mathsf{D}(\mathbb{Z})$ the objects $\mathbb{Z}_{p}$ and $\mathbb{Z}[p^{-1}]/\mathbb{Z}$ have the same derived endomorphism ring but generate different localizing subcategories. The slogan is: giving a proxy-small object $P$ is the same thing as giving a presentation of the compactly generated localizing subcategory $\operatorname{\mathrm{loc}}(P)$ as a quotient of $\mathsf{D}(\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(P,P))$ (see 4.10).
@@
Acknowledgements.
Briggs was partly supported by the European Union under Grant Agreement 10106455, and Iyengar was partly supported by National Science Foundation grant DMS-2001368. The authors thank John Greenlees for his comments, suggestions, and interest.
@@ 2. Proxy-smallness revisited
Let us fix a compactly generated stable $\infty$-category (i.e. an $\aleph_{0}$-presentable stable $\infty$-category) with homotopy category $\mathsf{K}$. The subcategory of compact objects in $\mathsf{K}$ is denoted $\mathsf{K}^{\mathrm{c}}$; it is, by definition, essentially small. Essentially $\mathsf{K}$ is a compactly generated triangulated category (see [14]*Remark 1.4.4.3) and we can speak of mapping spaces, homotopy (co)limits, and so on. In particular, given any two objects there is a spectrum of maps between them (see for instance Section 2.3 of [2] and in particular Definition 2.15) and hence any object has a ring spectrum of endomorphisms.
Little is lost if the reader prefers to think of instead fixing a differential graded, or model, category with compactly generated homotopy category. In fact, essentially all arguments take place at the level of homotopy categories. In the dg setting we would consider for any pair of objects the chain complex of maps between them and one could replace ring spectra by dg algebras.
Throughout by ‘localization’ we will mean a reflective localization of stable $\infty$-categories, i.e. an exact functor with a fully faithful right adjoint (what Lurie simply calls a localization in [16]*5.2.7.2 in the not necessarily stable setting). This is equivalent to a localization in the usual sense, i.e. given by inverting a suitable collection of maps in a homotopy coherent way, which moreover has a right adjoint (cf. [15] Proposition 04JL and the linked definitions). In the differential graded context this can be modelled by a Drinfeld quotient admitting a right adjoint.
Again, given the adjunction, this boils down to something which can be checked on the homotopy categories. The right adjoint is by definition fully faithful precisely when it induces weak equivalences of mapping spaces, and so given stability, exactly when it is fully faithful on the homotopy category. This can be rephrased as asking that the components of the counit are equivalences, i.e. isomorphisms in the homotopy category.
For a set of objects $\mathsf{X}$ of $\mathsf{K}$ we denote by $\operatorname{thick}(\mathsf{X})$ and $\operatorname{\mathrm{loc}}(\mathsf{X})$ the smallest triangulated subcategories containing $\mathsf{X}$ and closed under summands and arbitrary coproducts respectively.
@@
Throughout we regard $\mathbf{p}$ as a full subcategory of $\mathsf{K}$ and we use $i$ to denote the inclusion. Using the enhancement we may consider the ind-completion $\mathsf{D}(\mathbf{p})$ of $\operatorname{thick}(\mathbf{p})$ and this identifies with the full subcategory $\operatorname{\mathrm{loc}}(\mathbf{p})$ of $\mathsf{K}$ with $i^{*}$ the corresponding fully faithful left Kan extension, i.e. the inclusion of $\operatorname{\mathrm{loc}}(\mathbf{p})$. Details on ind-completion in general can be found in [16]*5.3 (see in particular 5.3.5.10 and 5.3.5.11) and see [14]*1.1.3.6 for the fact that this construction preserves stability.
@@
In fact one could, and in some cases should, go further: we could also replace $P$ by a set of objects; see 2.15. Everything works just as well if one takes this extra step. In our primary cases of interest we can get away with a single object $P$ and so we stick to this setting—replacing $P$ by a subcategory is essentially cosmetic.
@@
For the rest of this section $P$ is a proxy-small object in $\mathsf{K}$ with a set of proxies $\mathbf{p}$. We are primarily concerned with understanding the functor
$\Gamma=i_{*}i^{!}\cong\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(\mathbf{p},-)\otimes_{\mathbf{p}}\mathbf{p}.$ (2.7)
@@
Assume $P$ is proxy-small. The functor $(-)\otimes_{E}P\colon\mathsf{D}(E)\longrightarrow\operatorname{\mathrm{loc}}(P)$ is a localization. Moreover it preserves products.
@@
We verify that the functor $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(P,-)\colon\operatorname{\mathrm{loc}}(P)\longrightarrow\mathsf{D}(E)$ is fully faithful and right adjoint to $(-)\otimes_{E}P$. Our starting point is the observation that $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(P,-)$ induces an equivalence from $\operatorname{thick}(P)$ to $\mathsf{D}(E)^{\mathrm{c}}$ with inverse $(-)\otimes_{E}P$. In particular, $\mathbf{p}\subseteq\operatorname{thick}(P)$ embeds fully faithfully into $\mathsf{D}(E)$ which induces a colimit and compact preserving embedding $\lambda\colon\operatorname{\mathrm{loc}}(\mathbf{p})\longrightarrow\mathsf{D}(E)$. While the localizing subcategory $\operatorname{\mathrm{loc}}(P)$ must be taken in $\mathsf{K}$, we note that $\operatorname{\mathrm{loc}}(\mathbf{p})$ is unambiguous here: $\mathbf{p}$ consists of objects that are compact in both $\mathsf{D}(E)$ and $\mathsf{K}$, and so it does not matter where we form the localizing subcategory $\mathbf{p}$ generates. By Brown representability $\lambda$ has a right adjoint $\rho$. Of course, this is overkill: the right adjoint is just the obvious restriction functor (up to identifying $\operatorname{\mathrm{loc}}(\mathbf{p})$ and $\mathsf{D}(\mathbf{p})$).
@@
Consider $R=k[x]/(x^{2})$. In this case, $E=k[\theta]$ with $\theta$ in (homological) degree $-1$ and no higher homotopical structure (i.e. $E$ is formal). We have
@@
The remarks in the previous example extend to any artinian commutative local ring, and in fact it is more general still.
@@
Let $R$ be a coherent ring and $G$ a classical generator for $\mathsf{D}^{\mathrm{b}}(\operatorname{\mathsf{mod}}R)$, that is to say, $\operatorname{thick}(G)=\mathsf{D}^{\mathrm{b}}(\operatorname{\mathsf{mod}}R)$. Then $G$ is certainly proxy-small: $\operatorname{thick}(G)$ contains all perfect complexes and hence $\operatorname{\mathrm{loc}}(G)=\mathsf{D}(R)$, since the latter is generated by $R$.
Hence, letting $E$ denote the derived endomorphism ring of $G$, we see $\mathsf{D}(R)$ is a localization of $\mathsf{D}(E)$ and $\mathsf{D}(E)^{\mathrm{c}}\cong\mathsf{D}^{\mathrm{b}}(\operatorname{\mathsf{mod}}R)$. Thus $\mathsf{D}(E)\cong\mathsf{K}(\operatorname{\mathsf{Inj}}R)$ and this localization is the usual one, namely the functor inverting quasi-isomorphisms, up to equivalence.
We note that not all rings, not even commutative ones, have classical generators.
@@
Using Theorem 2.10 we can give an efficient proof of effective constructibility as in [7]*Theorem 4.10. The functor $\Gamma$ is the one from (2.7).
@@
$-\otimes_{\mathbb{Z}_{p}}P\colon\mathsf{D}(\mathbb{Z}_{p})\longrightarrow\mathsf{D}(\mathbb{Z})$
and so this functor is not fully faithful.
@@ 3. Detecting smashing localizations
We continue the notation of the previous sections, with the caveat that we no longer assume that $P$ is necessarily proxy-small.
If $P$ is proxy-small then the functor $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(P,-)\otimes_{E}P$ preserves coproducts, even though $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(P,-)$ does not if $P$ is not compact. It is then natural to ask what we can deduce about a general $P$ with the property that $\operatorname{\mathbf{R}Hom}_{\mathsf{K}}(P,-)\otimes_{E}P$ preserves coproducts.
@@
Since $\mathsf{K}$ is compactly generated $\operatorname{\mathrm{loc}}(P)$ is well-generated [19, Theorem 4.4.9] and so by Brown representability [19, Theorem 8.3.3] there is for all $X$ a localization triangle
@@ 4. Characterization of proxy-smallness
Remark 4.5.
In fact the localization corresponding to a proxy-small object also preserves products. This was shown in Theorem 2.10.
@@
One can also prove the preceding result by checking the definition of proxy-smallness directly using that $\mathsf{J}=\operatorname{\mathrm{loc}}(\pi(E))$, because $\pi$ is a localization, and applying Lemma 4.1.
@@
Here is another consequence of Theorem 4.3.
Corollary 4.11.
Let $R$ be a commutative noetherian ring, and $\{R\longrightarrow S_{\lambda}\}_{\lambda}$ a family of flat maps of rings such that the induced map $\sqcup_{\lambda}\operatorname{Spec}S_{\lambda}\longrightarrow\operatorname{Spec}R$ is onto. For an object $P$ in $\mathsf{D}^{\mathrm{b}}(\operatorname{\mathsf{mod}}R)$, if $S_{\lambda}\otimes_{R}P$ is proxy-small in $\mathsf{D}(S_{\lambda})$ for all $\lambda$, then $P$ is proxy-small.
Proof.
To begin with, we claim a map $f$ in $\mathsf{D}(R)$ is an isomorphism if $S_{\lambda}\otimes_{R}f$ is an isomorphism in $\mathsf{D}(S_{\lambda})$ for each $\lambda$.
Indeed, by considering the mapping cone of $f$, it suffices to verify that for an object $M\in\mathsf{D}(R)$, if $S_{\lambda}\otimes_{R}M=0$ in $\mathsf{D}(S_{\lambda})$ for all $\lambda$, then $M=0$ in $\mathsf{D}(R)$.
Fix $\mathfrak{p}$ in $\operatorname{Spec}R$ and pick an $S_{\lambda}$ and a $\mathfrak{q}\in\operatorname{Spec}S_{\lambda}$ such that $\mathfrak{q}\cap R=\mathfrak{p}$. Then, with $k(\mathfrak{p})$ and $k(\mathfrak{q})$ denoting the residue fields of $R_{\mathfrak{p}}$ and $S_{\mathfrak{q}}$, respectively, we get isomorphisms
$k(\mathfrak{q})\otimes_{S}(S_{\lambda}\otimes_{R}M)\cong k(\mathfrak{q})\otimes_{R}M\cong k(\mathfrak{q})\otimes_{k(\mathfrak{p})}(k(\mathfrak{p})\otimes_{R}M)\,.$
Thus we obtain that $k(\mathfrak{p})\otimes_{R}M=0$. Since $\mathfrak{p}$ was arbitrary, it follows that $M=0$, as claimed; see, for instance, [18, Lemma 2.12].
Set $E=\operatorname{\mathbf{R}Hom}_{R}(P,P)$. Since $R$ is noetherian and the $R$-module $H(P)$ is finitely generated, for any flat map $R\longrightarrow S$, and $M$ in $\mathsf{D}(R)$, the natural map is an isomorphism:
$S\otimes_{R}\operatorname{\mathbf{R}Hom}_{R}(P,M)\xrightarrow{\ \cong\ }\operatorname{\mathbf{R}Hom}_{S}(S\otimes_{R}P,S\otimes_{R}M)\,.$
We write $L^{R}_{P}(-)$ for the functor $\operatorname{\mathbf{R}Hom}_{R}(P,-)\otimes_{E}P$ on $\mathsf{D}(R)$. Given the isomorphism above, for any $M$ in $\mathsf{D}(R)$ and flat map $R\longrightarrow S$ one gets a natural isomorphism:
$S\otimes_{R}L^{R}_{P}(M)\longrightarrow L^{S}_{S\otimes_{R}P}(S\otimes_{R}M)\,.$
Now we are ready to verify that $P$ is proxy-small: Since the telescope conjecture holds for $\mathsf{D}(R)$, by [18, Corollary 3.4], it suffices to verify that for any family $\{M_{i}\}$ of $R$-complexes the natural map
$f\colon\bigoplus_{i}L^{R}_{P}(M_{i})\longrightarrow L^{R}_{P}(\bigoplus_{i}M_{i})$
is an isomorphism; see Corollary 4.6. Given the discussion above, it suffices to verify that $S_{\lambda}\otimes_{R}f$ is an isomorphism for each $\lambda$. Applying ($\ast$) one gets that
$S_{\lambda}\otimes_{R}f\colon\bigoplus_{i}L^{S_{\lambda}}_{S_{\lambda}\otimes_{R}P}(S_{\lambda}\otimes_{R}M_{i}))\longrightarrow L^{S_{\lambda}}(\bigoplus_{i}(S_{\lambda}\otimes_{R}M_{i}))\,.$
This is an isomorphism by Theorem 4.3 because $S_{\lambda}\otimes_{R}P$ is proxy-small, by hypothesis. ∎
It is illustrative to further specialise the preceding corollary; it recovers [13, Propositions 5.3 and 5.5].
Corollary 4.11.
Let $R$ be a commutative noetherian ring and $P$ in $\mathsf{D}^{\mathrm{b}}(\operatorname{\mathsf{mod}}R)$. The following conditions are equivalent:
  • (1)
    $P$ is proxy-small;
  • (2)
    $P_{\mathfrak{p}}$ is proxy-small in $\mathsf{D}(R_{\mathfrak{p}})$ for each $\mathfrak{p}\in\operatorname{Spec}R$;
  • (3)
    $\widehat{R}\otimes_{R}P$ is proxy-small in $\mathsf{D}(\widehat{R})$ where $\widehat{R}$ is the completion of $R$ with respect to an ideal in the Jacobson radical of $R$;
  • (4)
    $S\otimes_{R}P$ is proxy-small in $\mathsf{D}(S)$ for some faithfully flat map of rings $R\longrightarrow S$. ∎
Returning to the general context, when $\mathsf{K}$ has a compact generator each compactly generated localizing subcategory comes with a canonical presentation. Let $\mathbf{p}$ be a set of compact objects of $\mathsf{K}$ and set